# Friend or Foe? Harnessing Controllable Overfitting for Anomaly Detection

### v1

@@
Abstract
Overfitting has long been stigmatized as detrimental to model performance, especially in the context of anomaly detection. Our work challenges this conventional view by introducing a paradigm shift—recasting overfitting as a controllable and strategic mechanism for enhancing model discrimination capabilities. In this paper, we present Controllable Overfitting-based Anomaly Detection (COAD), a novel framework designed to leverage overfitting for optimized anomaly detection. We propose the Aberrance Retention Quotient (ARQ), a novel metric that systematically quantifies the extent of overfitting, enabling the identification of an optimal ”golden overfitting interval” $ARQ_{optimal}$. Within this interval, overfitting is leveraged to significantly amplify the model’s sensitivity to anomalous patterns, while preserving generalization to normal samples. Additionally, we present the Relative Anomaly Distribution Index (RADI), an innovative metric designed to complement AUROC-pixel by providing a more versatile and theoretically robust framework for assessing model performance. RADI leverages ARQ to track and evaluate how overfitting impacts anomaly detection, offering an integrated approach to understanding the relationship between overfitting dynamics and model efficacy. Our theoretical work also rigorously validates the use of Gaussian noise in pseudo-anomaly synthesis, providing the foundation for its broader applicability across diverse domains. Empirical evaluations demonstrate that our controllable overfitting method not only achieves State-Of-The-Art(SOTA) performance in both one-class and multi-class anomaly detection tasks but also redefines overfitting from a modeling challenge into a powerful tool for optimizing anomaly detection.
1 Introduction
In the traditional context of machine learning, overfitting has long been perceived as an undesirable phenomenon, just like  [3, 8, 10, 20], traditionally viewed as the result of a model excessively memorizing training data, leading to compromised generalization capabilities. However, we challenge this conventional wisdom by extending the concept beyond the domain of anomaly detection, positioning controllable overfitting as a versatile mechanism for enhancing model sensitivity across a wide range of applications. This leads us to the central question that underpins our work: can overfitting be a friend rather than a foe in the domain of anomaly detection?
@@ 1 Introduction
Figure 1: Each subfigure presents a sequence from left to right, displaying: the original image, the ground truth, the results from other methods, like RD [6], RD++[23], UniAD [28] and DiAD [11], and our results after overfitting enhancement based on existing frameworks with COAD. (a) illustrates a significant reduction in the false positive rate after applying our overfitting approach. (b) highlights the reduction in the false negative rate, demonstrating enhanced sensitivity towards true anomalies with our method. The consistent improvement across both metrics indicates the effectiveness of our methods in refining detection accuracy.
The conventional understanding of overfitting necessitates a significant paradigm shift. Historically stigmatized as inherently detrimental to machine learning models, overfitting, as our research demonstrates, can under specific controlled conditions be repurposed as a powerful mechanism for enhancing model performance. In scenarios where training data consists solely of normal samples, we show that allowing the model to comprehensively memorize these features enhances its sensitivity to deviations—namely, anomalies—during testing. This reframing positions overfitting not as an inherent flaw to be avoided, but as a strategic asset to be leveraged effectively.
Anomaly detection in visual domains is a crucial task that aims to identify patterns that do not conform to expected behavior. Traditionally, anomaly detection relies on the notion of learning the distribution of normal data and distinguishing anomalies by identifying outliers. Various methods have been proposed to achieve this, including reconstruction-based approaches [31, 15, 30], embedding-based models [25, 32, 13, 9], and anomaly synthesis technique [4, 28, 17, 23].
@@
Our theoretical framework and empirical validation jointly support this proposition. We demonstrate that, as the degree of overfitting increases, the score distribution for normal samples evolves accordingly, while the score distribution for anomalous samples remains stable. By introducing the Aberrance Retention Quotient, called ARQ, as an overfitting parameter, coupled with a noise loss term to finely control the degree of overfitting, we have identified an optimal ”Golden Overfitting Interval” $ARQ_{optimal}$ that maximizes detection efficacy without compromising model reliability. This strategic balance enables robust differentiation between normal and anomalous samples, providing a substantial performance boost.
We also introduce a novel metric, the Relative Anomaly Distribution Index, called RADI, to differentiate from traditional AUROC-pixel measures. Unlike AUROC-pixel, RADI is derived from the cumulative distribution function (CDF) of the normal and anomalous pixel prediction score distributions. While RADI provides an indirect reflection of detection performance and maintains a similar trend to AUROC-pixel, it also offers a more versatile framework for theoretical modeling and practical discussion of model capabilities. This distinction underscores its value as a complementary tool, adding a new layer of analytical depth to our approach.
Moreover, we substantiate the rationale for employing Gaussian noise in pseudo-anomaly detection modules. Through calculating the Total Variation Distance (TVD) between actual anomaly data and a Gaussian distribution across the entire dataset, we obtain a fair result. This outcome indicates a high level of similarity between true anomalies and a Gaussian distribution, thereby providing a theoretical validation for the common practice of using Gaussian noise as an effective pseudo-anomaly generator in previous works like GLASS [4] and SimpleNet [17]. This insight not only strengthens our methodology but also offers a theoretical foundation for broader applications in anomaly detection frameworks.
@@
  • We introduce the ARQ to precisely regulate the overfitting degree and the RADI as a complementary metric to AUROC-pixel, providing a more adaptable tool for mathematical modeling and facilitating Dual Control Mechanism with the use of COAD.
  • Our approach dismantles the demonization of overfitting, repurposing it as a generalizable module and achieving SOTA results in both one-class and multi-class anomaly detection tasks. We also provide a theoretical foundation for utilizing Gaussian noise as a preliminarily pseudo-anomaly generator.
2 Related Work
2.1 Overview of Anomaly Detection
@@
Reconstruction-based methods are founded on the hypothesis that models trained exclusively on normal data can successfully reconstruct normal instances while failing to do so for anomalies, thereby using reconstruction error as a measure of abnormality. Notable reconstruction-based approaches include Autoencoders (AE)[29, 33], Generative Adversarial Networks (GANs)[7, 15], and Reverse Distillation (RD)[6, 23]. Specific methods under these categories, such as DSR[31], RealNet[32], and DRAEM[30], attempt to learn the representation of normal data and generate accurate reconstructions for normal samples but tend to produce significant residuals for anomalous inputs.
@@
Recent developments have seen the use of embedding-based methods, besides reconstruction-based methods. These models leverage pre-trained networks to extract features from input images, aiming to separate anomalies from normal samples within the feature space. Anomalies are identified based on the distance between their embeddings and those of normal samples. Techniques such as Double-MMD RGP[25], RealNet[32], PyramidFlow[13], and CFlowAD[9] utilize this paradigm, some of which employ a memory bank to store representative features and apply distance metrics to detect anomalies. Embedding-based methods are particularly valuable in high-dimensional feature spaces, where the intricate relationships between different regions of the data can be better captured using a pre-trained model’s learned representations.
@@
Another line of research focuses on synthetic anomaly generation, a strategy that involves creating pseudo-anomalies to augment training data and improve detection performance. These synthetic anomalies are used to provide explicit anomaly labels in the training set, converting the unsupervised task into a form of supervised learning. Notable techniques include CutPaste [14], which pastes cut-out normal patches in different positions, and methods such as GLASS[4], UniAD[28], SimpleNet[17], RealNet[32],and DDPM [12], all of which employ Gaussian noise to generate synthetic anomalies at both feature and image levels. These approaches help in training the model to differentiate between normal and abnormal samples, thus improving its ability to detect true anomalies. By contrast, methods such as DRAEM[30] and DMDD [16] employ Berlin noise or other forms of noise, which have not demonstrated the same level of effectiveness in anomaly detection tasks. In subsequent sections, we will theoretically justify the use of Gaussian noise in these pseudo-anomaly generation modules, establishing its effectiveness and reasonableness as an anomaly synthesis strategy.
2.2 One-class vs. Multi-class
@@
One-class anomaly detection focuses on modeling the distribution of normal data to identify deviations. Methods such as RD [6] and RD++ [23] are foundational in this area. RD employs a pretrain model for feature extraction, while RD++ further enhances this process for improved robustness.
@@
Multi-class anomaly detection handles diverse object categories and intra-class variations. Notable approaches like UniAD [28] and HVQ-Trans [18] effectively generalize across multiple classes.
2.3 Non-Diffusion vs. Diffusion-Based
For the purpose of making our method easier for readers to understand and apply, we have also supplemented it with the diffusion-based framework.
@@
Non-diffusion methods, such as UniAD [28], have been effective for anomaly detection but often face limitations in preserving the semantic consistency of anomalous regions, especially in complex settings.
@@
Diffusion-based methods have recently gained traction due to their exceptional reconstruction capabilities, as evidenced by DiAD [11], which stands out among contemporary methods. DiAD employs a diffusion framework for generating anomalies while maintaining semantic information.
3 Method
3.1 Overview
(a) Training Stage
(b) Inference Stage
Figure 2: Comprehensive Framework for COAD.(a) Training Stage: In the traditional framework, the teacher model remains frozen throughout. The pseudo-anomaly generation and filtering components are optional, remaining unfrozen during the normal training stage but frozen during the controllable overfitting stage, under the guidance of the Dual Control Mechanism. The student model also remains unfrozen during normal training stage, with selective freezing employed during the controllable overfitting stages. When the Dual Control Mechanism fails, Freeze Command Signals are triggered, prompting the selective freezing of certain network layers in the traditional framework to effectively manage overfitting. (b) Inference Stage: The inference pipeline follows a similar procedure to the traditional framework, ensuring that the enhancements provided by COAD integrate seamlessly.
As  Fig. 2 shows, our works provides an overview of our proposed framework for enhancing anomaly detection through controllable overfitting. We introduce Controllable Overfitting-based Anomaly Detection, called COAD. We systematically integrate novel elements such as ARQ for quantifying overfitting and RADI as a complementary metric for AUROC-pixel. These components collectively facilitate the monitoring, control, and optimization of the overfitting process to maximize model sensitivity while preserving generalization capabilities.
3.2 Quantifying Overfitting with Aberrance Retention Quotient
The exploration of overfitting has largely been confined to its avoidance, and there has been a conspicuous lack of systematic studies on how to control or even beneficially leverage overfitting for enhancing model capabilities. Moreover, existing research lacks a quantitative metric that accurately reflects the extent of overfitting, which is critical for understanding its implications and finding the ”sweet spot” in model training. To fill this gap, we introduce the Aberrance Retention Quotient , called ARQ, a novel metric designed to quantify the degree of overfitting by capturing the model’s divergence from true data representation during training.
The ARQ is formally defined as follows:
$ARQ=\frac{\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})}{\sum_{i=1}^{N}y_{i}},$ (1)
@@ 3.2 Quantifying Overfitting with Aberrance Retention Quotient
ARQ is a core metric in our framework that quantifies the progression of overfitting during model training. By tracking ARQ, we identify the golden overfitting interval, where the model’s sensitivity to anomalous patterns is maximized without compromising generalization. Specifically, we denote the interval of optimal ARQ as:
$ARQ_{\text{optimal}}=[\theta-\delta,\theta+\delta],$ (2)
@@
3.3 Bridging Theoretical Modeling and Practical Anomaly Detection with Relative Anomaly Distribution Index
3.3.1 Probability Distribution Models of Normal and Anomalous Pixel Scores
To facilitate the mathematical treatment later, we made the following two assumptions, which were verified in subsequent experiments, please see  Sec. 4.
@@ 3.3.1 Probability Distribution Models of Normal and Anomalous Pixel Scores
$S_{n}\sim N(\mu_{n}(\theta),\sigma_{n}(\theta)^{2}),$ (3)
@@
$\sigma_{n}(\theta)=\sigma_{n0}e^{-k\theta},$ (4)
@@
$S_{a}\sim N(\mu_{a},\sigma_{a}^{2}),$ (5)
since anomalous pixels are not present during training, and the model cannot memorize them.
3.3.2 Introduction of RADI
After these assumptions, we introduce the Relative Anomaly Distribution Index, called RADI, a novel metric designed to bridge theoretical modeling and practical anomaly detection. Unlike traditional metrics such as AUROC, which are often used to evaluate model performance in anomaly detection tasks, RADI provides an indirect yet insightful measure of the model’s capacity to differentiate between normal and anomalous samples. This makes it particularly valuable for theoretical analyses, while also retaining practical utility in assessing model effectiveness.
The RADI quantifies the overlap between normal and anomalous score distributions, akin to the Wilcoxon rank-sum [24] and Mann-Whitney U tests [19]. RADI is calculated using the cumulative distribution function (CDF) of the score distributions:

### v2

@@
Abstract.
Overfitting has traditionally been viewed as detrimental to anomaly detection, where excessive generalization often limits models’ sensitivity to subtle anomalies. Our work challenges this conventional view by introducing Controllable Overfitting-based Anomaly Detection (COAD), a novel framework that strategically leverages overfitting to enhance anomaly discrimination capabilities. We propose the Aberrance Retention Quotient (ARQ), a novel metric that systematically quantifies the extent of overfitting, enabling the identification of an optimal ”golden overfitting interval” $ARQ_{optimal}$ wherein model sensitivity to anomalies is maximized without sacrificing generalization. To comprehensively capture how overfitting affects detection performance, we further propose the Relative Anomaly Distribution Index (RADI), a metric superior to traditional AUROC by explicitly modeling the separation between normal and anomalous score distributions. Theoretically, RADI leverages ARQ to track and evaluate how overfitting impacts anomaly detection, offering an integrated approach to understanding the relationship between overfitting dynamics and model efficacy. We also rigorously validate the statistical efficacy of Gaussian noise as pseudo-anomaly generators, reinforcing the method’s broad applicability. Empirical evaluations demonstrate that our controllable overfitting method achieves State-Of-The-Art(SOTA) performance in both one-class and multi-class anomaly detection tasks, thus redefining overfitting as a powerful strategy rather than a limitation.
Anomaly Detection, Controllable Overfitting, Aberrance Retention Quotient, Relative Anomaly Distribution Index, Gaussian Noise, Dual Control Mechanism
ccs: Computing methodologies Scene anomaly detection
1. Introduction
In the traditional context of machine learning, overfitting has long been perceived as an undesirable phenomenon, just like  (Bishop, 2006; Goodfellow et al., 2016; Hastie et al., 2004; Murphy, 2012), traditionally viewed as the result of a model excessively memorizing training data, leading to compromised generalization capabilities. However, we challenge this conventional wisdom by extending the concept beyond the domain of anomaly detection, positioning controllable overfitting as a mechanism for enhancing model sensitivity across a wide range of applications. This leads us to the central question that underpins our work: can overfitting be a friend rather than a foe in the domain of anomaly detection?
@@ 1. Introduction
Figure 1. Each subfigure presents a sequence from left to right, displaying: the original image, the ground truth, the results from other methods, like RD (Deng and Li, 2022), RD++(Tien et al., 2023), UniAD (You et al., 2022) and DiAD (He et al., 2023), and our results after overfitting enhancement based on existing frameworks with COAD. (a) illustrates a significant reduction in the false positive case after applying our overfitting approach. (b) highlights the reduction in the false negative case, demonstrating enhanced sensitivity towards true anomalies with our method. The consistent improvement across both metrics indicates the effectiveness of our methods in refining detection accuracy.
The conventional understanding of overfitting necessitates a significant paradigm shift. Historically stigmatized as inherently detrimental to machine learning models, overfitting, as our work demonstrates, can under specific controlled conditions be repurposed as a powerful mechanism for enhancing model performance. Intuitively, by enabling the model to intentionally and moderately memorize fine-grained patterns unique to normal samples, we amplify its sensitivity toward any deviations—precisely the subtle anomalies we aim to detect. This reframes overfitting from a phenomenon to avoid into a strategically leveraged resource.
Anomaly detection is a crucial task that aims to identify patterns that do not conform to expected behavior. Traditionally, anomaly detection relies on the notion of learning the distribution of normal data and distinguishing anomalies by identifying outliers. Various methods have been proposed, including reconstruction-based approaches (Zavrtanik et al., 2022; Liang et al., 2023; Zavrtanik et al., 2021), embedding-based models (Xiao et al., 2023; Zhang et al., 2024; Lei et al., 2023; Gudovskiy et al., 2021), and anomaly synthesis technique (Chen et al., 2024; You et al., 2022; Liu et al., 2023; Tien et al., 2023).
@@
Existing anomaly detection metrics such as AUROC-pixel predominantly evaluate performance at discrete thresholds, failing to systematically quantify how overfitting impacts model sensitivity across continuous distribution shifts. To fill this critical gap, we introduce the Aberrance Retention Quotient(ARQ), explicitly measuring the extent of overfitting by quantifying prediction deviations from normal data. Leveraging ARQ, we define a ”Golden Overfitting Interval” ($ARQ_{optimal}$), wherein controlled overfitting maximizes anomaly detection sensitivity without compromising model stability. Additionally, we propose the Relative Anomaly Distribution Index (RADI), an innovative metric derived directly from cumulative distribution functions of normal and anomalous predictions, capturing subtle changes in their overlap and thus providing a more informative assessment than AUROC alone.
Moreover, we substantiate the rationale for employing Gaussian noise in pseudo-anomaly detection modules. Through calculating the Total Variation Distance (TVD) between actual anomaly data and a Gaussian distribution across the entire dataset, we obtain a fair result. This outcome indicates a high level of similarity between true anomalies and a Gaussian distribution, thereby providing a theoretical validation for the common practice of using Gaussian noise as an effective pseudo-anomaly generator in previous works like GLASS (Chen et al., 2024) and SimpleNet (Liu et al., 2023). This insight not only strengthens our methodology but also offers a theoretical foundation for broader applications in anomaly detection frameworks.
@@
  • We introduce the ARQ to precisely regulate the overfitting degree and the RADI as a complementary metric to AUROC-pixel, providing a more adaptable tool for mathematical modeling and facilitating Dual Control Mechanism.
  • Our method achieves SOTA results in both one-class and multi-class anomaly detection tasks. We also provide a theoretical foundation for utilizing Gaussian noise as a preliminarily pseudo-anomaly generator.
2. Related Work
2.1. Overview of Anomaly Detection
@@
Reconstruction-based methods are founded on the hypothesis that models trained exclusively on normal data can successfully reconstruct normal instances while failing to do so for anomalies, thereby using reconstruction error as a measure of abnormality. Notable reconstruction-based approaches include Autoencoders (AE)(Zavrtanik, 2021; Zhou, 2022), Generative Adversarial Networks (GANs)(Goodfellow et al., 2014; Liang et al., 2023), and Reverse Distillation (RD)(Deng and Li, 2022; Tien et al., 2023). Specific methods under these categories, such as DSR(Zavrtanik et al., 2022), RealNet(Zhang et al., 2024), and DRAEM(Zavrtanik et al., 2021), attempt to learn the representation of normal data and generate accurate reconstructions for normal samples but tend to produce significant residuals for anomalous inputs.
@@
Recent developments have seen the use of embedding-based methods, besides reconstruction-based methods. These models leverage pre-trained networks to extract features from input images, aiming to separate anomalies from normal samples within the feature space. Anomalies are identified based on the distance between their embeddings and those of normal samples. Techniques such as Double-MMD RGP(Xiao et al., 2023), RealNet(Zhang et al., 2024), PyramidFlow(Lei et al., 2023), and CFlowAD(Gudovskiy et al., 2021) utilize this paradigm, some of which employ a memory bank to store features and apply distance metrics to detect anomalies. Embedding-based methods are particularly valuable in high-dimensional feature spaces, where the intricate relationships between different regions of the data can be better captured using a pre-trained model’s learned representations.
@@
Another line of research focuses on synthetic anomaly generation, a strategy that involves creating pseudo-anomalies to augment training data and improve detection performance. These synthetic anomalies are used to provide explicit anomaly labels in the training set, converting the unsupervised task into a form of supervised learning. Notable techniques include CutPaste (Li et al., 2021), which pastes cut-out normal patches in different positions, and methods such as GLASS(Chen et al., 2024), UniAD(You et al., 2022), SimpleNet(Liu et al., 2023), RealNet(Zhang et al., 2024),and DDPM (Ho et al., 2020), all of which employ Gaussian noise to generate synthetic anomalies at both feature and image levels. These approaches help in training the model to differentiate between normal and abnormal samples, thus improving its ability to detect true anomalies. By contrast, methods such as DRAEM(Zavrtanik et al., 2021) and DMDD (Liu et al., 2024) employ Berlin noise or other forms of noise, which have not demonstrated the same level of effectiveness in anomaly detection tasks. In subsequent sections, we will theoretically justify the use of Gaussian noise in these pseudo-anomaly generation modules, establishing its effectiveness and reasonableness as an anomaly synthesis strategy.
2.2. One-class vs. Multi-class
@@
One-class anomaly detection focuses on modeling the distribution of normal data to identify deviations. Methods such as RD (Deng and Li, 2022) and RD++ (Tien et al., 2023) are foundational in this area. RD employs a pretrain model for feature extraction, while RD++ further enhances this process for improved robustness.
@@
Multi-class anomaly detection handles diverse object categories and intra-class variations. Notable approaches like UniAD (You et al., 2022) and HVQ-Trans (Lu et al., 2023) effectively generalize across multiple classes.
2.3. Non-Diffusion vs. Diffusion-Based
For the purpose of making our method easier for readers to apply, we have supplemented it with the diffusion-based framework.
@@
Non-diffusion methods, such as UniAD (You et al., 2022) and GLAD (Yao et al., 2024), have been effective for anomaly detection but often face limitations in preserving the semantic consistency of anomalous regions, especially in complex settings.
@@
Diffusion-based methods have recently gained traction due to their exceptional reconstruction capabilities, as evidenced by DiAD (He et al., 2023), which stands out among contemporary methods. Besides, AnomalyDiffusion (Hu et al., 2024) and TransFusion (Fučka et al., 2024), which jointly learn to generate anomalies and discriminate them from normal data by use of Diffusion-based methods.
3. Method
3.1. Overview
Figure 2. Comprehensive Framework for COAD. In the Traditional Model, Teacher Model remains frozen throughout. The pseudo-anomaly Gaussian Noise is optional. The student model also remains unfrozen during normal training stage, with selective freezing employed during the controllable overfitting stages. When the Dual Control Mechanism fails, Freeze Command Signals are triggered, prompting the selective freezing of certain network layers and flagging the selected layer in the Traditional Model to effectively manage overfitting. The inference pipeline follows a similar procedure to the Traditional Model, ensuring that the enhancements provided by COAD integrate seamlessly.
As  fig. 2 shows, our work provides an overview of our proposed framework for enhancing anomaly detection through controllable overfitting. We introduce Controllable Overfitting-based Anomaly Detection, called COAD. We systematically integrate novel elements such as ARQ for quantifying overfitting and RADI as a complementary metric for AUROC-pixel. These components collectively facilitate the monitoring, control, and optimization of the overfitting process to maximize model sensitivity while preserving generalization capabilities. The detailed overall training and inference pipeline’s pseudo code can be seen in  algorithm 1 and  algorithm 3 in Supplementary Materials.
Given an input image $x\in\mathcal{X}$, our Teacher network $\mathcal{T}$, which is trained on a large-scale dataset, produces reference features: $y\;=\;\mathcal{T}(x)$. The Student network $\mathcal{S}$ then predicts $\hat{y}\;=\;\mathcal{S}(x)$. Finally, the discrepancy between $y$ and $\hat{y}$ determines the anomaly region. We get the anomaly region $M\;=\;y\;\oplus\;\hat{y}$.
[…]
3.2. Quantifying Overfitting with Aberrance Retention Quotient
The exploration of overfitting has largely been confined to its avoidance, yet systematic methods to quantify, control, and beneficially leverage overfitting remain largely unexplored. To fill this gap, we introduce the Aberrance Retention Quotient, called ARQ, a novel metric designed to quantify the degree of overfitting by capturing the model’s divergence from true data representation.
Specifically, the ARQ is formally defined as follows:
(1) $ARQ=\frac{\sum_{i=1}^{N}|\hat{y}_{i}-y_{i}|}{\sum_{i=1}^{N}y_{i}},$
@@ 3.2. Quantifying Overfitting with Aberrance Retention Quotient
ARQ is a core metric in our framework that quantifies the progression of overfitting during model training. Through continuous monitoring of ARQ, we establish an optimal interval golden overfitting interval, where the model’s sensitivity to anomalous patterns is maximized without compromising generalization. To objectively determine this interval, we employ a validation-based adaptive search strategy, optimizing ARQ to improve anomaly detection. Specifically, we denote the interval of optimal ARQ as:
(2) $ARQ_{\text{optimal}}=[\theta-\delta,\theta+\delta],$
@@
3.3. Bridging Theoretical Modeling and Practical Anomaly Detection with Relative Anomaly Distribution Index
3.3.1. Probability Distribution Models of Normal and Anomalous Pixel Scores
To facilitate the mathematical treatment later, we make the following two assumptions. We empirically verify these assumptions and find that normal distributions provide adequate approximations for prediction score distributions, a common phenomenon justified by central limit theorem and supported by empirical observations (refer to  fig. 5 and Supplementary Material).
@@ 3.3.1. Probability Distribution Models of Normal and Anomalous Pixel Scores
(3) $S_{n}\sim N(\mu_{n}(\theta),\sigma_{n}(\theta)^{2}),$
@@
(4) $\sigma_{n}(\theta)=\sigma_{n0}e^{-k\theta},$
@@
(5) $S_{a}\sim N(\mu_{a},\sigma_{a}^{2}),$
Since anomalous pixels represent unseen deviations that the model does not explicitly memorize during training, we assume their predicted score distribution remains relatively stable and unaffected by varying degrees of overfitting(refer to  fig. 5).
3.3.2. Introduction of RADI
After these assumptions, we introduce the Relative Anomaly Distribution Index, called RADI, a novel metric designed to bridge theoretical modeling and practical anomaly detection. Unlike threshold-dependent metrics such as AUROC, RADI directly measures the probabilistic separation between normal and anomalous distributions, providing more explicit insights into the model’s discriminative capability independent of specific thresholds. This makes it particularly valuable for theoretical analyses, while retaining practical utility in assessing model effectiveness.
The RADI quantifies the overlap between normal and anomalous score distributions, akin to the Wilcoxon rank-sum (Wilcoxon, 1992) and Mann-Whitney U tests (Mann and Whitney, 1947). RADI is calculated using the cumulative distribution function (CDF) of the score distributions: